Solve using elimination method. Label each system as consistent or inconsistent.
Thank you.
w + 8s = - 42
-9w - s = -48

Answer by checkley77(12844)

w + 8s = - 42 multiply by 9 & add.
-9w - s = -48
9w+72s=-378
--------------------
71s=-426
s=-426/71
s=-6 ans.
w+8*-6=-42
w-48=-42
w=-42+48
w=6 ans.
Proof:
-9*6-(-6)=-48
-54+6=-48
-48=-48